# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 5 Triangularity of transition matrices
The first part is [47, Proposition 6.1.9]. (See also [35, Proposition 18].) It implies that the multiplication map $\overline{BwB}\times^{B}\hat{S}\to G$ has connected fibres. Taking inverses in $G$, ($x\mapsto x^{-1}$), one concludes that a map like the multiplication map $\hat{S}\times^{B}\overline{BwB}\to G$ also has connected fibres. ∎
@@
The result also holds if ${\sf k}$ is a field of characteristic zero, but we will not need that.
Proposition 5.8.
Let $\lambda$ be a dominant weight.
  1. 1.
    $H^{0}(\overline{BwB}S,\mathcal{L}(\lambda))\to H^{0}({\overline{BwB}/B},\mathcal{L}(H^{0}(S,\mathcal{L}(\lambda)))$ is an isomorphism.
  2. 2.
    $H^{0}(\hat{S}\overline{BwB},\mathcal{L}(\lambda))\to H^{0}(S,\mathcal{L}(H^{0}({\overline{BwB}/B},\mathcal{L}(\lambda)))$ is an isomorphism.
  3. 3.
    $H^{i}({\overline{BwB}/B},\mathcal{L}(H^{0}(S,\mathcal{L}(\lambda)))=0$ for $i>0$.
  4. 4.
    $H^{i}(S,\mathcal{L}(H^{0}({\overline{BwB}/B},\mathcal{L}(\lambda)))=0$ for $i>0$.
@@
First assume ${\sf k}$ is a field of finite characteristic.
Proof of part 1. Consider the multiplication map $m:\overline{BwB}\times^{B}S\to\overline{BwB}S$. It has connected fibres, is separable, proper, has a target that is Frobenius split, so by [47, Proposition 6.1.6] we have $m_{*}\mathcal{O}_{\overline{BwB}\times^{B}S}=\mathcal{O}_{\overline{BwB}S}$. By the projection formula we have $m_{*}m^{*}\mathcal{L}(\lambda)=\mathcal{L}(\lambda)$. So $H^{0}(\overline{BwB}S,\mathcal{L}(\lambda))\cong H^{0}(\overline{BwB}\times^{B}S,m^{*}\mathcal{L}(\lambda))$. And $\overline{BwB}\times^{B}S\to\overline{BwB}$ is a Serre fibration with fibre $S$, so $H^{0}(\overline{BwB}\times^{B}S,m^{*}\mathcal{L}(\lambda))\cong H^{0}({\overline{BwB}/B},\mathcal{L}(H^{0}(S,\mathcal{L}(\lambda)))$.
Part 2 follows in a similar way.
Part 3 is basically the fact that $H^{0}(S,\mathcal{L}(\lambda))$ has relative Schubert filtration and is therefore ${\sf H}_{w}$ acyclic, cf. Lemma 5.3.
Part 4. By the previous part the result holds when $S$ is a Schubert variety. There are only finitely many possibilities for $S$, so we may assume the result for any union of Schubert varieties that is strictly contained in $S$. If $S$ is not a Schubert variety, write $S=S^{\prime}\cup S^{\prime\prime}$ where $S^{\prime}$, $S^{\prime\prime}$ are strictly smaller. Put $M=H^{0}({\overline{BwB}/B},\mathcal{L}(\lambda))$. Consider the Mayer–Vietoris sequence
$0\to H^{0}(S,\mathcal{L}(M))\to H^{0}(S^{\prime},\mathcal{L}(M))\oplus H^{0}(S^{\prime\prime},\mathcal{L}(M))\to H^{0}(S^{\prime}\cap S^{\prime\prime},\mathcal{L}(M))\to\cdots.$
By part 1 the map $H^{0}(S^{\prime\prime},\mathcal{L}(M))\to H^{0}(S^{\prime}\cap S^{\prime\prime},\mathcal{L}(M))$ may be identified with the surjective map $H^{0}(\overline{BwB}S^{\prime\prime},\mathcal{L}(\lambda))\to H^{0}(\overline{BwB}(S^{\prime}\cap S^{\prime\prime}),\mathcal{L}(\lambda))$. The result follows.
By the universal coefficient theorem all four parts go through when ${\sf k}={\mathbb{Z}}$ or when ${\sf k}$ is a field of characteristic zero [29, I Proposition 4.18, 5.8(2)].
@@ 5.1 The matrix $\beta_{vw}$
Recall that $\beta_{vw}=\langle[\mathcal{O}_{X^{w}}(-\partial X^{w})],[{\mathcal{P}}_{v}]\rangle$. We want to pair $[\mathcal{O}_{X^{w}}(-\partial X^{w})]$ with $[{\mathcal{P}}_{v}]=[\mathcal{L}(P(-e_{v}))]$. As $0\to\mathcal{O}_{X^{w}(-\partial X^{w})}\to\mathcal{O}_{X^{w}}\to\mathcal{O}_{\partial X^{w}}\to 0$ is exact, it suffices to compute $\langle[\mathcal{O}_{X^{w}}],[\mathcal{L}(P(-e_{v}))]\rangle$ minus $\langle[\mathcal{O}_{\partial X^{w}}],[\mathcal{L}(P(-e_{v}))]\rangle$.
Note that $X^{w}=\overline{Bww_{0}x_{-}}$.
And $P(-e_{v})=\Gamma(\overline{Bv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$.
We hope that $\langle[X^{w}(-\partial X^{w})],[{\mathcal{P}}_{v}]\rangle$ vanishes unless $vw_{0}\leq w$. But the propositions above show that $\langle[\mathcal{O}_{X^{w}}],[\mathcal{L}(P(-e_{v}))]\rangle$ is just the character of the module $\Gamma(\overline{Bww_{0}\star v^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$. Similarly $\langle[\mathcal{O}_{\partial X^{w}}],[\mathcal{L}(P(-e_{v}))]\rangle$ is just the character of the module $\Gamma(\partial^{w,v},\mathcal{L}(-w_{0}ve_{v}))$, where $\partial^{w,v}$ equals $(\partial\overline{Bww_{0}B})\;\overline{Bv^{-1}w_{0}B}x_{-}$. Both modules have a relative Schubert filtration and we can get a grip on them by inspecting the extremal weights. (The weights that have the same length as the dominant weight $ve_{v}$.)
@@
  • $vw_{0}=w$. So that is about $Q(\overline{Bw_{0}x_{-}},\partial^{vw_{0},v},\mathcal{L}(-w_{0}ve_{v}))$, the kernel of the surjective map from $\Gamma(\overline{Bw_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ to $\Gamma(\partial^{w,v},\mathcal{L}(-w_{0}ve_{v}))$. One checks it to be $Q(-ve_{v})={\sf k}_{-ve_{v}}$.
  • $vw_{0}\nleq w$. Then one sees that $Q(\overline{Bww_{0}\star v^{-1}w_{0}x_{-}},\partial^{w,v},\mathcal{L}(-w_{0}ve_{v}))$ vanishes.
@@ 5.2 The matrix $\alpha_{vw}$
Let $\lambda$ be an anti-dominant weight. And let $S$ be a $B^{+}$ invariant closed subset of $G/B^{+}$.
@@
    $H^{0}(\overline{B^{+}wB^{+}}S,\mathcal{L}^{+}(\lambda))\to H^{0}({\overline{B^{+}wB^{+}}/B^{+}},\mathcal{L}^{+}(H^{0}(S,\mathcal{L}^{+}(\lambda)))$ is an isomorphism.
@@
    $H^{i}({\overline{B^{+}wx_{+}}},\mathcal{L}^{+}(H^{0}(S,\mathcal{L}^{+}(\lambda)))=0$ for $i>0$. ∎
We now need to understand $\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(\phi_{0}^{*}(Q(e_{v}))))$, the kernel of the map $\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(P^{+}(w_{0}e_{v})))\to\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$, which is a surjective map by Lemma 5.1, Lemma 5.9.
@@
    $\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}})$ contains the union of the $\overline{B^{+}w_{0}v^{-1}w_{0}sx_{+}}$ with $s$ simple, $s\notin J$.
@@
    $ve_{v}$ is an extremal weight of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(P^{+}(w_{0}e_{v})))$, but not of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$.
@@
    If $uve_{v}\neq ve_{v}$, then $uve_{v}$ is an extremal weight of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v})))$.
@@
Part (1). The extremal weights of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(P^{+}(w_{0}e_{v})))$ are elements in the $W$ orbit of $w_{0}e_{v}$. The socle of $P^{+}(w_{0}e_{v})$ has weight $w_{0}e_{v}=w_{0}v^{-1}w_{0}w_{0}ve_{v}$ and $\ell(vw_{0})+\ell(w_{0}v^{-1}w_{0})=\ell(w_{0})$, so $(vw_{0})\star(w_{0}v^{-1}w_{0})=w_{0}$, and $w_{0}w_{0}ve_{v}$ is a weight of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(P^{+}(w_{0}ve_{v})))=\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(H^{0}(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}},\mathcal{L}^{+}(w_{0}ve_{v})))$. Now suppose $ve_{v}$ is an extremal weight of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$. Then $ve_{v}$ can be written as $y(w_{0}ve_{v})$ with $y\leq vw_{0}\star z$ for some $z<w_{0}v^{-1}w_{0}$. Replacing $z$ by a lesser element if necessary, we may assume $y\leq vw_{0}z$ for some $z<w_{0}v^{-1}w_{0}$. Put $u=w_{0}y$. Then $ve_{v}=yw_{0}ve_{v}=w_{0}uw_{0}ve_{v}$ with $u\geq w_{0}vw_{0}z$. But $uw_{0}ve_{v}=w_{0}ve_{v}$ implies $u\in W_{J}$, so $u\geq w_{0}vw_{0}z$ implies $w_{0}vw_{0}z\in W_{J}$. This contradicts the minimality of $w_{0}v^{-1}w_{0}$ in its coset $w_{0}v^{-1}w_{0}W_{J}$.
Part (2). Now consider a weight of the form $uve_{v}$ with $uve_{v}\neq ve_{v}$. Then $w_{0}uw_{0}w_{0}ve_{v}\neq w_{0}ve_{v}$ so $w_{0}uw_{0}\notin W_{J}$. There is a simple $s$, $s\notin J$, with $w_{0}uw_{0}\geq s$. Then $uw_{0}\leq w_{0}s=(vw_{0})\star(w_{0}v^{-1}w_{0}s)$ and $\overline{w_{0}v^{-1}w_{0}sx_{+}}$ is contained in $\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}})$ by Lemma 5.10. So $uw_{0}w_{0}ve_{v}$ is an extremal weight of $\Gamma(\overline{B^{+}vw_{0}x_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$. ∎
@@
Let $v,w\in W$ such that $w\nleq vw_{0}$. Every extremal weight of $\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(P^{+}(w_{0}e_{v})))$ is also an extremal weight of $\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v})))$.
@@
The extremal weights of $\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(P^{+}(w_{0}e_{v})))$ are of the form $uw_{0}ve_{v}$ with $u\leq w\star(w_{0}v^{-1}w_{0})$. Suppose $\ell(w)+\ell(w_{0}v^{-1}w_{0})=\ell(ww_{0}v^{-1}w_{0})$. Then $w\leq w_{0}w_{0}vw_{0}$, contrary to the assumption. So $\ell(w)+\ell(w_{0}v^{-1}w_{0})<\ell(ww_{0}v^{-1}w_{0})$ and we may chose $z<w_{0}v^{-1}w_{0}$ with $w\star(w_{0}v^{-1}w_{0})=w\star z$. As $u\leq w\star z$, $uw_{0}ve_{v}$ is an extremal weight of $\Gamma(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(\Gamma(\partial(\overline{B^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$. ∎
@@
6 ${\mathrm{ind}}$-vanishing for $B$-modules
In this section we draw conclusions about $B$-modules by combining the vanishing in Corollary 2.5 with Theorem 3.1 on triangular transition matrices.
6.1 Cohomological descent from $G/B$ to $B$

### v3

@@ 5 Triangularity of transition matrices
If ${\sf k}$ is a field of finite characteristic, then parts 1 and 2 follow from [48, Proposition 2.24, Theorem 1.9.(a)(ii)] by induction on $\ell(w)$. By the Universal coefficient Theorem 1.8 both parts go through when ${\sf k}={\mathbb{Z}}$ or when ${\sf k}$ is a field of characteristic zero.
Proof of part 3. By the previous parts the result holds when $S$ is a Schubert variety. There are only finitely many possibilities for $S$, so we may assume the result for any union of Schubert varieties that is strictly contained in $S$. If $S$ is not a Schubert variety, write $S=S^{\prime}\cup S^{\prime\prime}$ where $S^{\prime}$, $S^{\prime\prime}$ are strictly smaller. Put $X=\overline{{\mathbf{B}}w{\mathbf{B}}}/\mathbf{B}$ and $M=H^{0}(X,\mathcal{L}(\lambda))$. If ${\sf k}$ is a field of finite characteristic, then $S^{\prime}\cap S^{\prime\prime}$ is reduced by Ramanathan [10, Proposition 1.2.1, Chapter 2]. By [10, Corollary 1.6.6] it is then also reduced in characteristic zero. From the Mayer–Vietoris sequence
$0\to H^{0}(S,\mathcal{L}(M))\to H^{0}(S^{\prime},\mathcal{L}(M))\oplus H^{0}(S^{\prime\prime},\mathcal{L}(M))\to H^{0}(S^{\prime}\cap S^{\prime\prime},\mathcal{L}(M))\to\cdots$
we get
$\chi(S,\mathcal{L}(M))=\chi(S^{\prime},\mathcal{L}(M))+\chi(S^{\prime\prime},\mathcal{L}(M))-\chi(S^{\prime}\cap S^{\prime\prime},\mathcal{L}(M)).$ (5)
Similarly the Mayer–Vietoris sequence
$0\to H^{0}(\tilde{S}X,\mathcal{L}(\lambda))\to H^{0}(\tilde{S}^{\prime}X,\mathcal{L}(\lambda))\oplus H^{0}(\tilde{S}^{\prime\prime}X,\mathcal{L}(\lambda))\to H^{0}(\tilde{S}^{\prime}X\cap\tilde{S}^{\prime\prime}X,\mathcal{L}(\lambda))\to 0,$
gives
$\chi(\tilde{S}X,\mathcal{L}(\lambda))=\chi(\tilde{S}^{\prime}X,\mathcal{L}(\lambda))+\chi(\tilde{S}^{\prime\prime}X,\mathcal{L}(\lambda))-\chi(\tilde{S}^{\prime}X\cap\tilde{S}^{\prime\prime}X,\mathcal{L}(\lambda)).$ (6)
By Lemma 5.7 we know that $\tilde{S}^{\prime}X\cap\tilde{S}^{\prime\prime}X=(\tilde{S}^{\prime}\cap\tilde{S}^{\prime\prime})X$. As the right hand sides of equations (5), (6) agree, part 3 follows.
Instead of invoking [10, Corollary 1.6.6], we could choose $N>1$ so that all relevant $H^{i}(Y,{\mathcal{F}})$ are flat over ${\mathbb{Z}}[1/N]$ and use the Universal coefficient Theorem 1.8 to pass from the case of fields of finite characteristic to the case ${\sf k}={\mathbb{Z}}[1/N]$ and from there to the case where ${\sf k}$ is a field of characteristic zero. ∎
@@
We will not need that $H^{i}(S,\mathcal{L}(H^{0}({\overline{{\mathbf{B}}w{\mathbf{B}}}/{\mathbf{B}}},\mathcal{L}(\lambda)))$ vanishes for $i>0$.
Lemma 5.11.
Let $S$, $S^{\prime}$ be unions of Schubert varieties in $\mathbf{G}/{\mathbf{B}}$ with $S^{\prime}\subset S$ and let $\lambda$ be dominant. Put $M=\ker({{\sf H}}_{w}(H^{0}(S,\mathcal{L}(\lambda)))\to{{\sf H}}_{w}(H^{0}(S^{\prime},\mathcal{L}(\lambda))))$. Then $M$ has a relative Schubert filtration. The sections of the filtration are the $Q(\mu)$ with $\mu$ an extremal weight of ${{\sf H}}_{w}(H^{0}(S,\mathcal{L}(\lambda)))$ but not of ${{\sf H}}_{w}(H^{0}(S^{\prime},\mathcal{L}(\lambda)))$. The multiplicity of such $Q(\mu)$ is one, and $\mu\in W\lambda$.
@@
By Proposition 5.9 we may use Lemma 5.1 to conclude that $M$ has a relative Schubert filtration. To find out what the sections in the filtration are, it suffices to look at the characters. By Proposition 2.17 both ${{\sf H}}_{w}(H^{0}(S,\mathcal{L}(\lambda)))$ and ${{\sf H}}_{w}(H^{0}(S^{\prime},\mathcal{L}(\lambda)))$ have a character that is a ${\mathbb{Z}}$-linear combination of characters ${\mathrm{Char}}(Q(\nu))$ with $\nu\in W\lambda$. So $M$ also has such a character. But ${\mathrm{Char}}(Q(\nu))-e^{\nu}$ only involves weights $\tau$ with $(\tau,\tau)<(\nu,\nu)$. So by inspecting the extremal weights one gets the coefficients in the expansion of ${\mathrm{Char}}(M)$ in terms of the ${\mathrm{Char}}(Q(\nu))$. ∎
@@ 5.1 The matrix $\beta_{vw}$
Recall that $\beta_{vw}=\langle[\mathcal{O}_{{X}^{w}}(-\partial{X}^{w})],[{\mathcal{P}}_{v}]\rangle$. We want to pair $[\mathcal{O}_{{X}^{w}}(-\partial{X}^{w})]$ with $[{\mathcal{P}}_{v}]=[\mathcal{L}(P(-e_{v}))]$. As $0\to\mathcal{O}_{{X}^{w}(-\partial{X}^{w})}\to\mathcal{O}_{{X}^{w}}\to\mathcal{O}_{\partial{X}^{w}}\to 0$ is exact, it suffices to compute the difference between $\langle[\mathcal{O}_{{X}^{w}}],[\mathcal{L}(P(-e_{v}))]\rangle$ and $\langle[\mathcal{O}_{\partial{X}^{w}}],[\mathcal{L}(P(-e_{v}))]\rangle$.
Note that ${X}^{w}=\overline{{\mathbf{B}}ww_{0}x_{-}}$.
And $P(-e_{v})=\Gamma(\overline{{\mathbf{B}}v^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$.
We hope that $\langle[{X}^{w}(-\partial{X}^{w})],[{\mathcal{P}}_{v}]\rangle$ vanishes unless $vw_{0}\leq w$. But the propositions above show that $\langle[\mathcal{O}_{{X}^{w}}],[\mathcal{L}(P(-e_{v}))]\rangle$ is just the character of the module $\Gamma(\overline{{\mathbf{B}(}ww_{0}\star v^{-1}w_{0})x_{-}},\mathcal{L}(-w_{0}ve_{v}))$. Similarly $\langle[\mathcal{O}_{\partial{X}^{w}}],[\mathcal{L}(P(-e_{v}))]\rangle$ is just the character of the module $\Gamma(\partial^{w,v},\mathcal{L}(-w_{0}ve_{v}))$, where $\partial^{w,v}$ equals $(\partial\overline{{\mathbf{B}}ww_{0}{\mathbf{B}}})\;\overline{{\mathbf{B}}v^{-1}w_{0}{\mathbf{B}}}x_{-}$. Both modules have a relative Schubert filtration and we can get a grip on them by inspecting the extremal weights. (The weights that have the same length as the dominant weight $ve_{v}$.)
@@
  • $vw_{0}=w$. So that is about the kernel of the surjective map from $\Gamma(\overline{{\mathbf{B}}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ to $\Gamma(\partial^{w,v},\mathcal{L}(-w_{0}ve_{v}))$. One checks it to be $Q(-ve_{v})={\sf k}_{-ve_{v}}$.
  • $vw_{0}\nleq w$. Then one sees that the kernel of the surjective map from $\Gamma(\overline{{\mathbf{B}}(ww_{0}\star v^{-1}w_{0})x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ to $\Gamma(\partial^{w,v},\mathcal{L}(-w_{0}ve_{v}))$ vanishes.
@@ 5.2 The matrix $\alpha_{vw}$
Let $\lambda$ be an anti-dominant weight. And let $S$ be a ${\mathbf{B}}^{+}$ invariant closed subset of $\mathbf{G}/{\mathbf{B}}^{+}$.
@@
    $H^{0}(\overline{{\mathbf{B}}^{+}w{\mathbf{B}}^{+}}S,\mathcal{L}^{+}(\lambda))\to H^{0}({\overline{{\mathbf{B}}^{+}w{\mathbf{B}}^{+}}/{\mathbf{B}}^{+}},\mathcal{L}^{+}(H^{0}(S,\mathcal{L}^{+}(\lambda)))$ is an isomorphism.
@@
    $H^{i}({\overline{{\mathbf{B}}^{+}wx_{+}}},\mathcal{L}^{+}(H^{0}(S,\mathcal{L}^{+}(\lambda)))=0$ for $i>0$. ∎
By Lemma 5.3 we have
$\alpha_{vw}=\langle[\mathcal{O}_{{X}_{w}}],[{\mathcal{Q}}_{v}]\rangle=\Gamma(\overline{{\mathbf{B}}^{+}wx_{+}},\mathcal{L}^{+}(\phi_{0}^{*}(Q(e_{v}))))$. Now $\phi_{0}^{*}(Q(e_{v})))$ has $\mathbf{B}^{+}$-socle of weight $w_{0}e_{v}$ and it is therefore the kernel of the surjection $\Gamma(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}},\mathcal{L}^{+}(w_{0}ve_{v})){\twoheadrightarrow}\Gamma(\partial(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v})).$ Combining with Proposition 5.12 we see $\alpha_{vw}$ is the character of the kernel of the surjective map from $\Gamma(\overline{({\mathbf{B}}^{+}w{\mathbf{B}}^{+}})(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$ to $\Gamma(\overline{({\mathbf{B}}^{+}w{\mathbf{B}}^{+}})\partial(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$.
@@
    $\partial(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}})$ contains the union of the $\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}sx_{+}}$ with $s$ simple, $s\notin J$.
@@
    $ve_{v}$ is an extremal weight of $\Gamma(\overline{({\mathbf{B}}^{+}vw_{0}{\mathbf{B}}^{+}})(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$, but not of $\Gamma((\overline{{\mathbf{B}}^{+}vw_{0}\mathbf{B}^{+}})\partial(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$.
@@
    If $uve_{v}\neq ve_{v}$, then $uve_{v}$ is an extremal weight of $\Gamma(\overline{({\mathbf{B}}^{+}vw_{0}\mathbf{B}^{+}})\partial(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$.
@@
Part (1). The extremal weights of
$M:=\Gamma(\overline{({\mathbf{B}}^{+}vw_{0}{\mathbf{B}}^{+}})(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$
are elements in the $W$ orbit of $w_{0}e_{v}$. We have $\ell(vw_{0})+\ell(w_{0}v^{-1}w_{0})=\ell(w_{0})$, so $(vw_{0})\star(w_{0}v^{-1}w_{0})=w_{0}$, and $w_{0}w_{0}ve_{v}$ is one of the weights of $M$. Now suppose $ve_{v}$ is an extremal weight of $\Gamma((\overline{{\mathbf{B}}^{+}vw_{0}\mathbf{B}^{+}})\partial(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$. Then $ve_{v}$ can be written as $y(w_{0}ve_{v})$ with $y\leq vw_{0}\star z$ for some $z<w_{0}v^{-1}w_{0}$. Replacing $z$ by a lesser element if necessary, we may assume $y\leq vw_{0}z$ for some $z<w_{0}v^{-1}w_{0}$. Put $u=w_{0}y$. Then $ve_{v}=yw_{0}ve_{v}=w_{0}uw_{0}ve_{v}$ with $u\geq w_{0}vw_{0}z$. But $uw_{0}ve_{v}=w_{0}ve_{v}$ implies $u\in W_{J}$, so $u\geq w_{0}vw_{0}z$ implies $w_{0}vw_{0}z\in W_{J}$. This contradicts the minimality of $w_{0}v^{-1}w_{0}$ in its coset $w_{0}v^{-1}w_{0}W_{J}$.
Part (2). Now consider a weight of the form $uve_{v}$ with $uve_{v}\neq ve_{v}$. Then $w_{0}uw_{0}w_{0}ve_{v}\neq w_{0}ve_{v}$ so $w_{0}uw_{0}\notin W_{J}$. There is a simple $s$, $s\notin J$, with $w_{0}uw_{0}\geq s$. Then $uw_{0}\leq w_{0}s=(vw_{0})\star(w_{0}v^{-1}w_{0}s)$ and $\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}sx_{+}}$ is contained in $\partial(\overline{\mathbf{B}^{+}w_{0}v^{-1}w_{0}x_{+}})$ by Lemma 5.13. So $uw_{0}w_{0}ve_{v}$ is an extremal weight of $\Gamma((\overline{{\mathbf{B}}^{+}vw_{0}\mathbf{B}^{+}})\partial(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$. ∎
@@
Let $v,w\in W$ such that $w\nleq vw_{0}$. Every extremal weight of $\Gamma(\overline{({\mathbf{B}}^{+}w{\mathbf{B}}^{+}})(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$ is also extremal weight of $\Gamma(\overline{({\mathbf{B}}^{+}w{\mathbf{B}}^{+}})\partial(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$.
@@
The extremal weights of $\Gamma(\overline{({\mathbf{B}}^{+}w{\mathbf{B}}^{+}})(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$ are of the form $uw_{0}ve_{v}$ with $u\leq w\star(w_{0}v^{-1}w_{0})$. Suppose $\ell(w)+\ell(w_{0}v^{-1}w_{0})=\ell(ww_{0}v^{-1}w_{0})$. Then $w\leq w_{0}w_{0}vw_{0}$, contrary to the assumption. So $\ell(w)+\ell(w_{0}v^{-1}w_{0})<\ell(ww_{0}v^{-1}w_{0})$ and we may chose $z<w_{0}v^{-1}w_{0}$ with $w\star(w_{0}v^{-1}w_{0})=w\star z$. As $u\leq w\star z$, we see that $uw_{0}ve_{v}$ is an extremal weight of $\Gamma(\overline{({\mathbf{B}}^{+}w{\mathbf{B}}^{+}})\partial(\overline{{\mathbf{B}}^{+}w_{0}v^{-1}w_{0}x_{+}}),\mathcal{L}^{+}(w_{0}ve_{v}))))$
@@
6 ${\mathrm{ind}}$-vanishing for $\mathbf{B}$-modules
In this section we draw conclusions about $\mathbf{B}$-modules by combining the vanishing in Corollary 2.21 with Theorem 4.1 on triangular transition matrices.
6.1 Cohomological descent from $\mathbf{G}/\mathbf{B}$ to $\mathbf{B}$